This is a cross section of two nested spheres. The outer sphere is a thin conductor of radius M with a uniform charge distrib5.) (2pts) Determine the capacitance between this arrangement of spheres.
Answer #1

1) Net charge =-2 +2 + 一3 In the fig tos nested sphores chout $e disq for LJ (regione) €. 4002= P (493) E = fer for Jerem $ E as= E. 4m² = Px200 > Ex​​​​​​

Please upvote if it helps5) Capacitance b/w arrangement Sehat C= 9 kach - (A - -